OpenStudy (anonymous):

I WILL BECOME A FAN! (9^-6)^3=9

OpenStudy (anonymous):

let me think for a sec

OpenStudy (anonymous):

Ok! :)

OpenStudy (anonymous):

that equation is false the real answer is 6.6624633053756668558384673044919e-18

OpenStudy (anonymous):

@UltimateWitherLord is correct Your equation looks like this : \[\left(\begin{matrix}1 \\ 9^6\end{matrix}\right)^3 =9\] So adding and multiplying those are incorrect because if you do, Your answer would be way larger then 9 So the problem is false.

OpenStudy (anonymous):

Exactly? But i cant find the answer

OpenStudy (anonymous):

put your question in the calculator exactly how you wrote it out or do it on paper

OpenStudy (anonymous):

Break your equation down \[9^6\] = 9 x 9 x 9 x 9 x 9 x 9 And \[1^3\] = 1 x 1 x 1 Add and you will find that it is false

OpenStudy (anonymous):

So there must not be an answer for it?

OpenStudy (anonymous):

You have two answers for that problem : 1. It's False 2. The total answer for the problem would be 6.6624633053756668558384673044919-18 What does your problem give you for answers?

OpenStudy (anonymous):

It doesnt give me any answers it says put in the answer.

OpenStudy (espex):

Then put "false" because \(9^{-6}\neq 9^{-3}\)

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

I agree with @eSpeX